Probability

16.A researcher wishes to examine whether floor area of a house (SQFT) has a positive impact on a housing sales in an area of a city. To check this, a researcher collected a sample of 345 observations, and ran a simple linear regression Price 36050.74 40.38 SQFT (se) (16294.25) (11.39) r2 0.61 Which of the following statements is TRUE? A) B) C) D) The null hypothesis is rejected at 10% but not 5 % level of significance The null hypothesis is rejected at both 5% and 10% level of significance The degrees of freedom for the test are 343 Both B and C are correct. 17.Which of the following statements best describe the principle of least squares for c alculating the regression line? A) Choose the line that minimises the value of r2 B) Choose the line that maximises the sum of squares of the deviations the C) Choose the line that minimises the sum of squares of the deviation of the D) Choose the line that maximises the sum of squares of the deviations the predicted values of Y from the mean of the dependent variable observed values of Y from the predicted values of Y predicted values of Y from the mean of the dependent variable 18. Which of the following is NOT TRUE about the Students t distribution? A) It is used to construct confidence intervals for the population mean when B) It has more area in the tails and less in the centre than does the normal C) As the number of degrees of freedom increases, the t distribution D) It is bell shaped and symmetrical the population standard deviation is known distribution approaches the normal distribution
Answer #1

(16)

t = b/se = 40.38/11.39 = 3.5452

df = 345 - 2 = 343

p- value = 0.0004

Option D

(18)

Option C

(19)

Option A